# Opponent Shaping for Antibody Development

### v6

@@ Abstract
Anti-viral therapies are typically designed to target only the current strains of a virus. Game theoretically, this corresponds to a short-sighted, or myopic, response. However, therapy-induced selective pressures act on viruses to drive the emergence of mutated strains, against which initial therapies have reduced efficacy. Building on a computational model of binding between antibodies and viral antigens (the Absolut! framework), we design and implement a genetic simulation of viral evolutionary escape. Crucially, this allows our antibody optimisation algorithm to consider and influence the entire escape curve of the virus, i.e. to guide (or “shape”) the viral evolution. This is inspired by opponent shaping which, in general-sum learning, accounts for the adaptation of the co-player rather than playing a myopic best response. Hence we call the optimised antibodies shapers. Within our simulations, we demonstrate that our shapers target both current and simulated future viral variants, outperforming the antibodies chosen in a myopic way. Furthermore, we show that shapers exert specific evolutionary pressure on the virus compared to myopic antibodies. Altogether, shapers modify the evolutionary trajectories of viral strains and minimise the viral escape compared to their myopic counterparts. While this is a simplified model, we hope that our proposed paradigm will facilitate the discovery of better long-lived vaccines and antibody therapies in the future, enabled by rapid advancements in the capabilities of simulation tools. Our code is available at https://github.com/olakalisz/antibody-shapers.
Keywords Machine Learning  $\cdot$ Game Theory  $\cdot$ Antibody Design  $\cdot$ Opponent Shaping  $\cdot$ Computational Biology
@@ 1 Introduction
Designing effective therapies to fight off viral pathogens is crucial. However, traditional design approaches only target the current variant of a virus. Although this myopic design approach may yield therapies with high initial efficacy, it fails to account for viral adaptation, leaving treatments vulnerable to becoming ineffective over time [1, 2, 3, 4, 5]. The COVID-19 pandemic starkly illustrated the challenges of adaptive viruses. While the rapid development of vaccines was a remarkable achievement, concerns quickly arose about their long-term efficacy against new emerging COVID variants [6, 7]. For example, the B.1.351 variant demonstrated that the vaccine loses its effectiveness against new strains [8]. This underscores the need for approaches that consider both the current and future efficacy of a designed therapy.
The therapies we design inevitably exert a degree of influence over viral evolution, as the viruses adapt in response to selective pressures imposed by said therapies. Recognizing this, our work aims to develop a method for designing therapies that can influence viral evolution and steer it towards less dangerous variants. To achieve this we utilise principles from opponent shaping [9]. Opponent shaping is a multi-agent reinforcement learning framework that allows agents to anticipate and influence the future strategies of other agents in their environment. This approach, exemplified by methods such as Learning with Opponent-Learning Awareness (LOLA) [9] and Model-Free Opponent Shaping (M-FOS) [10], encourages agents to consider not only their immediate rewards but also the consequences of their actions on their opponents’ future behaviour after a learning update.
Building on these principles, our work focuses specifically on antibody design, framing the interaction between antibodies and surface viral antigens as a two-player zero-sum game. In this game, the antibody’s payoff is primarily determined by its binding strength to the virus, whilst the virus has the opposite payoff. Although our framework can use any binding model (or indeed any sequence-to-function map), in this work we use our GPU-accelerated implementation of the Absolut! framework [11] to approximate the binding strength of protein-protein interactions. We use this game to model viral escape - the process via which mutations in a virus allow it to successfully evade a host’s immune system [12]. In the language of game theory, this is the virus adapting to the current antibody by repeatedly finding an approximate best response, decreasing the virus’s binding strength to the antibody. Using this viral escape model we optimise antibodies to be effective over the entire process of viral evolution, we refer to these as shapers. This is in contrast to only optimising for binding to the initial virus, which results in myopic antibodies.
Importantly, our simulations reveal that myopic antibodies are suboptimal when considering a longer time horizon of viral escape. Shaper antibodies not only outperform their myopic counterparts in long-term efficacy but also exert a shaping influence on viral evolutionary trajectories, effectively constraining the virus’s ability to escape antibody binding. Our study also explores the trade-offs between the effectiveness of shaper antibodies and the computational resources required for their optimisation, providing insights into the scalability and potential for practical deployment of our method. Finally, we present an analysis of the key features that distinguish shapers from myopic antibodies.
In summary, this work introduces a novel antibody design philosophy and implementation framework, leveraging opponent-shaping principles to improve the long-term efficacy of antibodies against evolving viruses. While our results provide a promising proof of concept, they are based on simplified models of binding and viral escape, limiting direct real-world application. However, as more sophisticated methods for modelling protein interactions and predicting viral evolution emerge, our framework has the potential to significantly impact future antiviral therapy design.
2 Results
@@
Figure 1: Computational Implementation of Opponent Shaping for Antibody Design. a An overview of our implementation. In the Antibody Optimisation Loop (i.e. the outer loop), we optimise the antibody to perform well against current and future virus variants; thus influencing the viral evolution. We approximate the future variants through our Simulated Viral Escape via Evolution (i.e. the inner loop) where the viruses evolve to escape from the current antibody over a given horizon length. b The payoffs of the antibody and virus. Red arrows indicate binding interactions that players aim to minimise, while green arrows represent those they aim to maximise. The antibody optimises for binding to the virus while avoiding its anti-target. In this zero-sum game, the antibody’s optimisation indirectly counters the virus’s binding to its target, see Equation 1. c Efficient JAX [13] implementation of the binding calculation uses binding poses generated by Absolut! [11] and the Miyazawa-Jernigan energy potential matrix [14].
2.1 Computational Implementation of Opponent Shaping for Antibody Design
To implement opponent shaping for antibody design, we develop a computational framework that models the virus-antibody interaction and optimises antibodies for long-term efficacy. Our approach comprises three main components: defining the virus-antibody game with the player’s payoffs, simulating viral escape, and optimising antibody shapers. We cover each of these components separately in the following three sections.
2.1.1 Virus-Antibody Game
@@
We formalise the interaction between antibodies and viruses as a two-player zero-sum game. In this game, two players – the virus and the antibody – play a game where one player’s gain is the other’s loss. The game is defined by the set of actions available to each player and their respective payoffs. The players’ actions are represented by their amino acid sequences. The sequences are of an antigen protein for the virus and a fragment of a hypervariable region of the heavy chain for the antibody, specifically the third complementarity-determining region (CDRH3), which is the most important part of the antibody for defining its specificity and affinity [15].
@@
The payoff structure is designed to capture the biological incentives of both players: the antibody aims to bind strongly to the virus while avoiding binding to host proteins (an anti-target), whereas the virus seeks to evade antibody binding while maintaining its ability to bind to host cell receptors (a binding target). Mathematically, we define the antibody’s payoff $R_{a}$ as:
$R_{a}(v,a)=B(v,a)-B(t^{-}_{a},a)-B(v,t^{+}_{v})$ (1)
@@
where $B(v,a)$ represents the binding strength between the virus $v$ and antibody $a$, $t^{-}_{a}$ is the antibody’s anti-target, and $t^{+}_{v}$ is the virus’s binding target. The virus’s payoff $R_{v}$ is simply the negative of the antibody’s payoff: ${R_{v}(v,a)=-R_{a}(v,a)}$, see Figure 1b. This formulation also ensures that neither player can adopt an overly simplistic strategy: the antibody can’t become universally “sticky” due to the anti-target penalty, and the virus can’t become entirely inert without losing its ability to infect host cells. This game structure forms the foundation for our method, see Methods section 4.1 for more details.
To evaluate the binding strength $B(\cdot,\cdot)$ in our setup, we adapt the Absolut! framework [11]. Absolut! is a simplified simulator for antibody-antigen binding. Unlike sequence-based machine learning models [16, 17, 18, 19] that cannot generalise to novel viral mutations, or computationally intensive molecular dynamics simulations [20], Absolut! offers a fast approach that generalises to both viral and antibody mutations, which is necessary for our application. However, using Absolut! sacrifices accuracy, a trade-off we discuss more explicitly in Section 3.
Specifically, the Absolut! framework operates by discretising the antigen structure, and evaluating the interaction between the antigen and the CDRH3 region of the antibody’s heavy chain - the most variable portion of the antibody [15]. It enumerates all possible binding poses of the CDRH3 to the discretised antigen and computes their binding energy using the Miyazawa–Jernigan energy potential [14]. Absolut! outputs the binding energy $E$ of the lowest energy pose as the binding energy of a query antibody-antigen complex. The binding strength is high when the binding energy is low. Thus we define our binding function as $B(\cdot,\cdot)=-E(\cdot,\cdot)$, see Methods Section 4.4 for more details.

### v8

# ADIOS: Antibody Development via Opponent Shaping

@@ Abstract
Anti-viral therapies are typically designed to target only the current strains of a virus, a myopic response. However, therapy-induced selective pressures drive the emergence of new viral strains, against which the original myopic therapies are no longer effective. This evolutionary response presents an opportunity: our therapies could both defend against and actively influence viral evolution. This motivates our method ADIOS: Antibody Development vIa Opponent Shaping. ADIOS is a meta-learning framework where the process of antibody therapy design, the outer loop, accounts for the virus’s adaptive response, the inner loop. With ADIOS, antibodies are not only robust against potential future variants, they also influence, i.e., shape, which future variants emerge. In line with the opponent shaping literature, we refer to our optimised antibodies as shapers. To demonstrate the value of ADIOS, we build a viral evolution simulator using the Absolut! framework, in which shapers successfully target both current and future viral variants, outperforming myopic antibodies. Furthermore, we show that shapers modify the distribution over viral evolutionary trajectories to result in weaker variants. We believe that our ADIOS paradigm will facilitate the discovery of long-lived vaccines and antibody therapies while also generalising to other domains. Specifically, domains such as antimicrobial resistance, cancer treatment, and others with evolutionarily adaptive opponents. Our code is available at https://github.com/olakalisz/adios.
Machine Learning, Game Theory, Antibody Design, Opponent Shaping, Computational Biology
@@ 1 Introduction
Designing effective therapies to fight off viral pathogens is crucial for limiting their devastating social and economic costs (Nandi & Shet, ; Orenstein & Ahmed, 2017; Samsudin et al., 2024; Faramarzi et al., 2024). However, traditional design approaches only target the current variant of a virus. Although this myopic design approach may yield therapies with high initial efficacy, it fails to account for viral adaptation, leaving treatments vulnerable to becoming ineffective over time (Weisblum et al., 2020; Doud et al., 2018; Lee et al., 2019; Dingens et al., 2019; Greaney et al., 2021).
Figure 1: Myopic Therapies Become Ineffective Over Time. Initial virus (variant A) evolves in response to the evolutionary pressures induced by therapies, resulting in new variants. In case of traditional myopic therapies (top), the new emerging variants (variant B) are often therapy resistant. In contrast, ADIOS designs shapers or shaper therapies (bottom) which remain effective and steer the viral evolution towards less harmful variants (variant C).
The COVID-19 pandemic starkly illustrated the challenges of adaptive viruses. While the rapid development of vaccines was a remarkable achievement, concerns quickly arose about their long-term efficacy against new emerging COVID variants (Carabelli et al., 2023; Hu et al., 2021). For example, the B.1.351 variant demonstrated that the vaccine loses its effectiveness against new strains (Madhi et al., 2021). This underscores the need for approaches that consider both the current and future efficacy of a designed therapy.
The virus inevitably adapts in response to selective pressures imposed by our therapies, i.e., we influence the viral evolution (Chéron et al., 2016; Meijers et al., 2022). Our work turns this influence in our favour, designing therapies that steer the virus towards less dangerous variants, see Figure 1.
@@
Figure 2: Main Components of ADIOS. a The ADIOS framework. In the Antibody Optimisation Loop (i.e., outer loop), we optimise the antibody to perform well against current and future virus variants; thus influencing the viral evolution. We approximate the future variants through our Simulated Viral Escape via Evolution (i.e., inner loop) where the viruses evolve to escape from the current antibody over a given horizon length. b The payoffs of the antibody and virus. Red arrows indicate binding interactions that players aim to minimise, while green arrows represent those they aim to maximise. The antibody optimises for binding to the virus while avoiding its anti-target. In this zero-sum game, the antibody’s optimisation indirectly counters the virus’s binding to its target, see Equation 1. c Binding simulator. Our JAX (Bradbury et al., 2018) implementation of the binding calculation uses binding poses generated by Absolut! (Robert et al., 2022) and the Miyazawa-Jernigan energy potential matrix (Miyazawa & Jernigan, 1999).
To achieve this we utilise principles from opponent shaping (Foerster et al., 2018), a multi-agent reinforcement learning framework that allows agents to both anticipate and influence the future policies of other agents in their environment. This approach, exemplified by methods such as Learning with Opponent-Learning Awareness (Foerster et al., 2018) and Model-Free Opponent Shaping (Lu et al., 2022), allows agents to consider not only their current performance but also the consequences of their actions on their opponents’ future behaviour.
Building on these principles, we introduce ADIOS: Antibody Development vIa Opponent Shaping. Antibodies are immune system proteins that bind to pathogens such as viruses. While naturally produced by the body, it is also possible to design and synthetically produce antibodies as therapies. ADIOS frames the interaction between antibodies and viruses as a two-player zero-sum game. In this game, the antibody’s payoff is primarily determined by its binding strength to the virus, while the virus has the opposite payoff (Figure 2b). Although our framework can use any binding model in principle, in this work we build on the Absolut! framework (Robert et al., 2022) to estimate the binding strength of protein-protein interactions. To improve computational efficiency, we reimplement parts of Absolut! in JAX (Bradbury et al., 2018), allowing GPU acceleration and a 10,000-fold speedup over the original implementation (Figure 2c).
We use this game to model viral escape - the process through which mutations allow a virus to evade a host’s immune system (Lucas et al., 2001). Following a meta-learning approach, ADIOS implements two nested optimisation loops, an inner loop and an outer loop (Figure 2a). In the inner loop, we simulate viral escape via evolution, where the virus adapts to the current antibody by repeatedly finding approximate best responses that decrease binding strength. The outer loop uses a genetic algorithm to optimise the antibodies to be effective across viral evolutionary trajectories, resulting in antibodies we call shapers. This is in contrast to only optimising for binding to the initial virus, which results in myopic antibodies.
Importantly, our simulations show that shapers not only outperform myopic antibodies in long-term efficacy but also demonstrate the ability to shape viral evolution. Moreover, shapers guide viruses toward variants that are more susceptible to binding to a broad spectrum of antibodies, not just the shaper that induced the given viral evolution, providing insights into the scalability and potential for practical deployment of ADIOS. Our study also explores the trade-offs between the effectiveness of shapers and the computational resources required for their optimisation. Finally, we present an explainability analysis of the key features that distinguish shapers from myopic antibodies.
Our key contributions include:
  • ADIOS: A framework that brings opponent shaping to antibody design to address viral escape.
  • A GPU-accelerated JAX implementation of the binding simulator Absolut!, achieving a 10,000x speedup.
  • An open-source instantiation of ADIOS applied to antibody design for both the dengue virus and three other viruses using our JAX implementation.
  • Empirical results showing ADIOS-optimised shapers both significantly outperform myopic antibodies by limiting long-horizon viral escape and guide viral evolution towards variants that can be more easily targeted.
  • Analysis of computational trade-offs in shaping horizons, providing practical guidance for deploying ADIOS in compute-constrained settings, e.g. due to more realistic binding simulators.
  • Interpretability analysis into how antibody shapers influence viral escape, which could, in principle, provide inspiration to antibody designers.
While our results provide a promising proof of concept, they are based on simplified models of binding and viral escape. However, we believe that as more sophisticated simulators emerge, the ADIOS framework has the potential to significantly impact future antiviral therapy design.
2 Related Work
Antibody Design:
Antibodies are essential components of the immune system that bind to unique identifiers (antigens) present on pathogens, including viruses, to identify and neutralise them. While natural antibodies emerge through an immune response, it is also possible to design antibodies for use as therapies. Recent work has made significant progress in computational antibody design (Cutting et al., 2024; Zambaldi et al., 2024; Bennett et al., 2024). The common approaches to antibody design utilise energy-based antibody optimisation methods (Li et al., 2014; Adolf-Bryfogle et al., 2018; Pereira et al., 2024), sequence-based language models (Liu et al., 2020; Saka et al., 2021) or structure-based approaches relying on GNNs (Jin et al., 2022) and diffusion models (Martinkus et al., 2024).
In contrast to these works, we are not interested in generating better antibody design methods immediately, but rather in how we should make new methods in the future to account for our effect on evolving viruses.
Predicting Viral Escape:
Recent machine learning methods have demonstrated success in predicting future viral strains (Shanker et al., 2024; Wang et al., 2023; Nie et al., 2025). EVEscape (Thadani et al., 2023) decomposes the likelihood of a mutation into three parts: maintaining fitness, accessibility to antibodies, and disrupting binding, demonstrating success through retrospective identification of COVID variants. Han et al. (2023) take a different approach by modelling viral evolution through simulated fitness landscapes. Unlike these methods, ADIOS models the antibody influence on viral evolution, enabling both the simulation of viral escape trajectories and the optimisation of antibodies to minimise viral escape.
3 Background
Antibody Binding Simulators:
In our setting, the interaction between antibodies and viruses is characterised by their binding strength $B(\cdot,\cdot)$ a measure of how strongly the two “attach” to each other through molecular forces. Molecular dynamics simulations offer high accuracy but are computationally intensive (Hollingsworth & Dror, 2018). Sequence-based ML models (Mason et al., 2021; Lim et al., 2022; Ruffolo et al., 2023; Yan Huang et al., 2022) provide faster alternatives but struggle to generalise beyond their training distribution, making them unsuitable for exploring novel viral mutations. To evaluate $B(\cdot,\cdot)$ we use the Absolut! framework (Robert et al., 2022). Absolut! offers a balance between speed and generalisation by modelling binding through discretised protein structures. It focuses on the CDRH3 region of the antibody, the most variable portion that primarily determines binding specificity (VanDyk & Meek, 1992). For each antibody-antigen pair, Absolut! enumerates possible binding poses and computes their energy using the Miyazawa-Jernigan potential (Miyazawa & Jernigan, 1999). The binding strength $B(\cdot,\cdot)$ is then defined as the negative of the lowest binding energy, see Figure 2c and Appendix D for details.
Opponent Shaping:
A multi-agent reinforcement learning framework which allows agents to anticipate and influence the future policies of other agents in their environment. Learning with Opponent-Learning Awareness (LOLA) (Foerster et al., 2018) introduced this concept by having LOLA agents optimise against anticipated opponent updates rather than static opponent policies. They achieved this through an augmented value function that accounts for the opponent’s learning step.
Meta-learning is a set of methods for optimising a learning process itself, “learning to learn”. In multi-agent systems, this concept extends to learning about and influencing how other agents learn. Model-Free Opponent Shaping (M-FOS) (Lu et al., 2022) showcases this idea by using gradient-free optimisation to learn meta-policies that accomplish long-horizon opponent shaping. Our approach follows a similar principle, using evolutionary optimisation to shape viral escape trajectories.
4 Method
ADIOS frames antibody design as a two-player game between an antibody shaper agent and a naive virus agent, building on principles from opponent shaping (Figure 2). We present our method in three parts:
In Section 4.1, we introduce the virus-antibody game, defining the action spaces and payoffs for both players. Section 4.2 describes our simulated viral evolution process, modelling how viruses evolve to escape a given antibody. Finally, Section 4.3 presents our antibody optimisation approach, which optimises the antibody shapers in a way that accounts for future viral mutations and learns to influence viral evolution away from escape.
Algorithm 1 $\text{Ev}(\hat{v},a)$: Simulated Viral Escape (Inner Loop)
  Input: virus $\hat{v}$, antibody $a$, horizon $H$, population size $P$, inverse temperature $\beta$
  Output: Sampled viral trajectory $\bm{\hat{v}}=[\hat{v}^{0},\hat{v}^{1},\dots,\hat{v}^{H}]$
  $\hat{v}_{0}\leftarrow\hat{v}$
  for $i=0$ to $H-1$ do
     for $k=1$ to $P$ do
        $v^{i}_{k}\leftarrow\hat{v}^{i}\oplus\textit{Mutation}$
     end for
     $p(v)\leftarrow\mathbb{P}(v=v^{i}_{k})\propto\exp(\beta R_{\text{v}}(v^{i}_{k},a))$ // See Eq. 1
     $\hat{v}^{i+1}\sim p(v)$
  end for
  $\bm{\hat{v}}\leftarrow[\hat{v}_{0},\dots,\hat{v}_{H}]$
  return $\bm{\hat{v}}$
4.1 Virus-Antibody Game
@@
We formalise the interaction between antibodies and viruses as a two-player zero-sum game. In this game, two players – the virus and the antibody – play a game where one player’s gain is the other’s loss. The game is defined by the set of actions available to each player and their respective payoffs. The players’ actions are represented by their amino acid sequences. The sequences are of an antigen protein for the virus and a fragment of a hypervariable region of the heavy chain for the antibody.
We define the set of 20 amino acids as $\mathbb{A}$. Let $N_{v}$ be the virus sequence length and $N_{a}$ be the antibody sequence length. So an action of the virus is $v\in\mathbb{A}^{N_{v}}$, and an action of the antibody is $a\in\mathbb{A}^{N_{a}}$. Let $B:\mathbb{A}^{N_{v}}\times\mathbb{A}^{N_{a}}\rightarrow\mathbb{R}$ be our binding function, which measures the strength of the binding between the antibody and the virus with increasing values corresponding to stronger binding [^1] This is opposite to binding energies, which are smaller for stronger binding..
@@
The payoff structure is designed to capture the biological incentives of both players: the antibody aims to bind strongly to the virus while avoiding binding to human proteins (an anti-target), whereas the virus seeks to evade antibody binding while maintaining its ability to bind to host cell receptors (a binding target). Mathematically, we define the antibody’s payoff $R_{a}$ as:
$R_{a}(v,a)=B(v,a)-B(t^{-}_{a},a)-B(v,t^{+}_{v})\vspace{-0.2cm}$ (1)
@@
where $B(v,a)$ represents the binding strength between the virus $v$ and antibody $a$, $t^{-}_{a}$ is the antibody’s anti-target, and $t^{+}_{v}$ is the virus’s binding target. The virus’s payoff $R_{v}$ is simply the negative of the antibody’s payoff: ${R_{v}(v,a)=-R_{a}(v,a)}$, see Figure 2b. This formulation also ensures that neither player can adopt an overly simplistic strategy: the virus can’t become entirely inert without losing its ability to infect host cells, and the antibody can’t become universally “sticky” without binding to the human protein, a “false positive”, potentially causing the immune system to attack the human body. We give the full Markov Decision Process (MDP) definition $\mathcal{M}=\bigl{\langle}\mathcal{S},\,\mathcal{A}^{\mathrm{v}},\,\mathcal{A}^{\mathrm{a}},\,P,\,R,\,\mu\bigr{\rangle}$ in Appendix E.
Algorithm 2 Antibody Optimisation (Outer Loop)
  Input: antibody $\hat{a}$,virus $\hat{v}$, horizon $H$, population size $P_{a}$, steps $N$
  Output: Trajectory of antibodies $\bm{\hat{a}}=[\hat{a}^{0},\hat{a}^{1},\dots,\hat{a}^{N}]$
  $\hat{a}_{0}\leftarrow\hat{a}$
  for $i=0$ to $N-1$ do
     $a^{i}_{1}\leftarrow\hat{a}^{i}$
     for $k=2$ to $P_{a}$ do
        $a^{i}_{k}\leftarrow\hat{a}^{i}\oplus\textit{Point Mutation}$
     end for
     $\hat{a}^{i+1}\leftarrow\arg\max_{k}\mathbb{E}\left[F_{\hat{v}}^{H}(a^{i}_{k})\right]$ // See Eq. 2
  end for
  $\bm{\hat{a}}\leftarrow[\hat{a}_{0},\dots,\hat{a}_{N}]$
  return $\bm{\hat{a}}$
4.2 Simulated Viral Escape via Evolution
We model the viral escape as a virus $\hat{v}$ naively evolving for $H$ steps in response to some fixed antibody $a$. The simulated viral escape via evolution, see Figure 2a and Algorithm 1, is defined as follows. Given a starting virus $\hat{v}$, the fixed antibody $a$ induces a distribution $\text{Ev}(\hat{v},a)$ over sequences of viruses $\bm{\hat{v}}=[\hat{v}^{0},\hat{v}^{1},\hat{v}^{2},\dots\hat{v}^{H}]$, where $\hat{v}^{0}=\hat{v}$ and $H$ is the chosen horizon length. We write $\bm{\hat{v}}\sim\text{Ev}(\hat{v},a)$ to denote this relationship.
We define the process of generating the escape trajectories inductively. In generation $i$, we have a virus $\hat{v}^{i}$. We generate a population of viruses $v^{i}_{1},v^{i}_{2}\dots v^{i}_{P}$ by duplicating $\hat{v}^{i}$ $P$ times, then randomly applying mutations such that on average there is one amino acid mutation per viral sequence:
$v^{i}_{k}=\hat{v}^{i}\oplus\textit{Mutation}\vspace{-0.2cm}$
With duplicates in the population being considered distinct, so that the likelihood of a particular variant increases with the number of duplicates. Furthermore, $\beta$ is a constant which reflects how random the selection process is, with $\beta\rightarrow\infty$ reflecting deterministic max-fitness selection. After $H$ generations, a full escape trajectory $\bm{\hat{v}}=[\hat{v}^{0},\hat{v}^{1},\hat{v}^{2},\dots,\hat{v}^{H}]$ has been generated and the simulated viral escape process ends.
4.3 Antibody Optimisation
$F^{H}_{\hat{v}}(a)=\mathbb{E}_{\bm{\hat{{v}}}\sim\text{Ev}(\hat{v},a)}\left[\frac{1}{H+1}\sum^{H}_{i=0}R_{a}(\hat{v}^{i},a)\right]$ (2)
Note that if $H=0$ this fitness defaults to a naive antibody payoff that ignores viral escape, i.e., $F^{0}_{\hat{v}}(a)=R_{a}(\hat{v},a)$. We refer to this as the myopic objective.
To optimise both shapers and myopic antibodies, we employ Monte Carlo simulations to estimate the antibody fitness, combined with an evolutionary optimisation algorithm. We refer to this process as the antibody optimisation loop, see Figure 2a and Algorithm 2. In meta-learning terms, this is the outer loop or the meta-loop, contrasting to the inner loop, which is the simulated viral escape via evolution.
Given a starting antibody $\hat{a}$, a starting virus $\hat{v}$ and a viral escape horizon $H$, the antibody optimisation process generates a trajectory of antibodies $\bm{\hat{a}}=[\hat{a}^{0},\hat{a}^{1},\hat{a}^{2},\dots,\hat{a}^{N}]$, where $N$ is the number of antibody optimisation steps (i.e., meta-steps). In the trajectory, $\hat{a}^{0}=\hat{a}$ is the starting antibody and $\hat{a}_{N}$ is the final optimised antibody. This optimisation could, in principle, start from any antibody, but for simplicity we opt to start from purely random antibodies, meaning $\hat{a}$ is random. In most of our experiments $N=30$.
At the start of antibody optimisation step $i$, we have an antibody $\hat{a}^{i}$. We first generate a population of $P_{a}$ antibodies $[a^{i}_{1},a^{i}_{2}\dots a^{i}_{P_{a}}]$ by taking both the antibody $\hat{a}^{i}$ and $P_{a}-1$ copies of it, with each copy having exactly a single random mutation in the amino acid sequence of $\hat{a}^{i}$. For our experiments, $P_{a}=40$.
We then sample their fitness values $F^{H}_{\hat{v}}(a^{i}_{j})$ with a fixed number $\eta$ of Monte Carlo roll-outs, i.e., we sample $\eta$ independent viral escape trajectories, each with horizon $H$ viral escape steps. We found $\eta=5$ to be sufficient. Finally, we select $\hat{a}^{i+1}$ to be the best-performing antibody:
$\hat{a}^{i+1}=\arg\max_{k}\mathbb{E}\left[F_{\hat{v}}^{H}(a^{i}_{k})\right]$
Once the final optimised antibody $\hat{a}^{N}$ is generated, a full optimisation trajectory is complete, $\bm{\hat{a}}=[\hat{a}^{0},\hat{a}^{1},\hat{a}^{2},\dots,\hat{a}^{N}]$, and the antibody optimisation process finishes.
5 Experimental Setup
5.1 Absolut! Speedup